In January of 20X1, the Phillips Company purchased a patent at a cost of $100,000. In addition, $10,000 in legal fees were paid to acquire the patent. The company estimated a 10-year useful life for the patent and uses the straight-line amortization method for intangible assets. In 20X3, Phillips spent $25,000 in legal fees for an unsuccessful defense of the patent. The amount charged to income (expense and loss) in 20X3 related to the patent should be:

Multiple Choice

  • $103,000

  • $36,000

  • $113,000

  • All of these answer choices are incorrect.
Answer #1

Amortization of patent per year = (100,000+10,000)/10 = 11,000 per year

Unamortized patents in 20x3 = 110,000 - (11,000*2 years) = 88,000

Legal fees for unsuccessful defense = 25,000

Amount charged to income = 88,000 +25,000

= 113,000